Assume: Demand Curve: QD = 80 – 10P; and Supply Curve: QS = 10P 7. Given the information derived above, identify on the graph consumer surplus and producer surplus for each situation as well as deadweight loss, if any. b. Government imposes a minimum price of $6.00 Calculate and assess (describe the impact) of the following: 1. Consumer Surplus 2. Producer Surplus 3. Deadweight Loss 4. Total Surplus 5. Government Revenue 6. Is the market operating efficiently: Yes or No. Why? INCLUDE A GRAPH
Answer #1

(a) DD: 0 = so-lop. ss: Qs = 109 0o = os 80-10P = 10p 80 = 10P + 10p 80= 2op Q= loxy P=4 = 40 SS 9 VM DD 80 cs = — 18-42(40)min8=6 40 CS = 1 18-62 (20) = 4x2x20. = 20 PS = (6-2) (200+ IX (2-07 (20) = 4X20 + 2x2x20 = 80+ 20 = 100 DWL I16-27.140-20) =